rssLink RSS for all categories
 
icon_red
icon_green
icon_red
icon_red
icon_blue
icon_green
icon_green
icon_red
icon_red
icon_green
icon_orange
icon_green
icon_green
icon_green
icon_red
icon_blue
icon_red
icon_orange
icon_red
icon_red
icon_red
icon_red
icon_red
icon_red
icon_red
icon_red
icon_red
icon_orange
icon_green
 

FS#14461 — bhs-25b-ast

Attached to Project— Network
Maintenance
Beauharnois → BHS-1
CLOSED
100%
This morning (24/03/2017), we had to swap several optics on bhs-25b-ast due to massive logs by these optics.
The swap started at 13:49 GMT+1 and finish at 14:11 GMT+1.
We were late on the invervention due to a link which that did't go UP.

When we swap the second optic on bhs-25b-ast, we loss the equipement between 13:49 and 14:55. A BGP session was not properly configured.

Some may have been felt between 13:49 and 14:55 GMT+1.

All optics has been successfully swapped. They no longer generate bad logs.
Date:  Friday, 24 March 2017, 14:37PM
Reason for closing:  Done